A rectangle has an area of 27 squared inches. The height is three times the base. What is the perimeter of the rectangle?
: A rectangle has an area of 27 squared inches. The height is three times the base. What is the perimeter of the rectangle?

Answer by checkley77(3848)

+=plus
-=minus
*=multiply
/=divide
^=power.
-----------------------
h*b=27
3b*b=27
3b^2=27
b^2=27/3
b^2=9
b=sqrt9
b=3 in. is the base.
3*3=9 in. for the height.
Proof:
3*7=27
----------------------
2*3+2*9=6+18=24 in. is the perimeter.